I have a question about my homework so it says Some friends eat out and each pays for one-fifth of a pizza, one-half of a plate of fries, and one $2 soda. Each person paid a total of $10. How much does a plate of fries cost?
Answer by macston(5194)

.
P=cost of pizza; F=cost of fries; S=cost of soda
T=total cost per person
.
T=0.2P+0.5F+S
T-0.2P=0.5F+S
(T-0.2P-S)=0.5F
2(T-0.2P-S)=F
2($10-0.2P-$2)=F
2($8-0.2P)=F
$16-0.4P=F This is as far as we can go with information given.